OpenStudy (anonymous):

I have to evaluate the expression and type it out as an integer or simplified fraction. I will type it out in just a sec.............

OpenStudy (anonymous):

The equation looks like this only with the 18 and 17 and 19 and 18 on each line being low and equal\[1-\frac{ _{18^{P}_{17}} }{_{19^{P}_{18}} ? }\] Does anyone have any idea as to what this is? Or how to solve it? Or what the answer is? I need to solve it by simplifying it.

OpenStudy (anonymous):

(It doesn't have a question mark in the original problem by the way

ganeshie8 (ganeshie8):

\[\large 1-\frac{ ^{18}P_{17} }{^{19}P_{18}} \]

ganeshie8 (ganeshie8):

like this ?

OpenStudy (anonymous):

So like 18 - pick - 17?

OpenStudy (anonymous):

Yes like that @ganeshie8 . :-D

OpenStudy (anonymous):

Well no wait it is more like the P is the exponent of them.

OpenStudy (anonymous):

\[\huge 1-\frac{ _{18}P _{17} }{ _{19}P _{18} }\]

OpenStudy (anonymous):

YES. Like that @iambatman . :-D

Miracrown (miracrown):

Is this problem something to do with the permutation function? @ShortStaticBurst

OpenStudy (anonymous):

\[\large _{18}P _{17} = \frac{ 18! }{ (18-17)! }\] \[_{19}P _{18} = \frac{ 19! }{ (19-18)! }\] Do you know what permutation is?

Miracrown (miracrown):

Do you know if this expression is in the context of another problem?

OpenStudy (anonymous):

I know what permutation is somewhat. :-) But not entirely :-(

OpenStudy (anonymous):

I did like a little tutorial for it not too long ago..you can use it as a reference.

OpenStudy (anonymous):

This expression is not in the context of another problem. It asks you to work it by evaluating the system and type either an integer or a simplified fraction as the answer

Miracrown (miracrown):

if this is a permutation notation, we have the following: